# Machine-precision energy conservative quadrature hyperreduction of Lagrangian hydrodynamics

### v1

@@ 2.2 Temporal discretization
with $\bar{\bm{v}}_{k+1/2}=(\bm{v}_{k}+\bm{v}_{k+1})/2$. To ensure stability of the explicit timestepping scheme we employ the timestep control algorithm described in [72].
@@ 3 Reduced model
The FEM coefficient vectors of the reduced basis functions are represented by the orthonormal basis matrices $\bm{\Phi}_{v}\in\mathbb{R}^{N_{v}\times n_{v}}$, $\bm{\Phi}_{e}\in\mathbb{R}^{N_{e}\times n_{e}}$ and $\bm{\Phi}_{x}\in\mathbb{R}^{N_{v}\times n_{x}}$ where each column holds the FEM coefficients of one basis function. In this work we compute the reduced basis matrices from simulation data of the full model using the POD method [19]. To decide how many POD basis functions to use to form each basis matrix we employ the POD energy ratio
@@
where $\sigma_{i}\geq 0$ denotes the $i$th largest singular value of the employed data matrix consisting of $N_{snap}$ solution snapshots, $e_{\sigma}\geq 0$ a chosen threshold, and $n_{k}$ the size of the corresponding reduced state vector (velocity $v$, energy $e$ or position $x$). Using the introduced reduced basis matrices we rewrite (3.1) in terms of FEM coefficient vectors
@@
with the appropriate initial condition at $t=0$. The above corresponds to using the reduced basis functions as the trial functions and the FEM basis functions as the test functions. Next we present the derivation of two alternative reduced systems of equations for the reduced state vector $\hat{\bm{w}}=(\hat{\bm{v}},\hat{\bm{e}},\hat{\bm{x}})\in\mathbb{R}^{n}$, $n=n_{v}+n_{e}+n_{x}$. The reason for introducing two alternative systems is that both systems will be employed by the hyperreduction methods we introduce in the following sections.
Starting from (3.4) the first reduced system is derived by inverting the mass matrices and transpose multiplying by the reduced basis matrices
@@
Although the size of the reduced systems (3.5) and (3.7) has been successfully reduced to $n\ll N$, the nonlinearity of the force functions in the right hand side of the equations means that they still depend on the full state vector $\tilde{\bm{w}}\in\mathbb{R}^{N}$. As a result, evaluating the force functions requires lifting the reduced state $\hat{\bm{w}}\in\mathbb{R}^{n}$ to $\tilde{\bm{w}}$ at every timestep of the integration. The cost of this lifting operation dominates the computational cost of the simulation and prevents its acceleration. To address this issue a hyperreduction method is employed to derive a hyperreduced model that can be simulated using the reduced state $\hat{\bm{w}}$ without requiring lifting to the full state $\tilde{\bm{w}}$.
In this work we use a hyperreduction method called the empirical quadrature procedure (EQP) that uses reduced quadrature rules to evaluate the nonlinear force functions with reduced cost [57, 58]. Quadrature methods such as EQP are naturally well suited to FEM spatial discretization procedures such as the one employed in this work. The present work can be seen as a continuation of [33, 34] where interpolation methods were used to derive hyperreduced models for Lagrangian hydrodynamics problems.
@@ 3.1 Empirical quadrature procedure
In this section we describe the EQP hyperreduction method using the model equations (3.5). The calculation of the force vectors $\bm{F}_{v}^{\psi}$ and $\bm{F}_{e}^{\psi}$ of (3.5) requires the spatial integrations
@@
where $\bm{g}_{v}^{\psi}$ and $\bm{g}_{e}^{\psi}$ are the integrands taking values in $\mathbb{R}^{n_{v}}$ and $\mathbb{R}^{n_{e}}$ respectively with components
@@
These integrations can be approximated using numerical quadrature rules employing the underlying FEM basis functions [72, 77]. Let $\{x^{v}_{j},\rho^{v}_{j}\}_{j=0}^{J_{v}-1}$ and $\{x^{e}_{j},\rho^{e}_{j}\}_{j=0}^{J_{e}-1}$ denote the full quadrature rules ($x_{j}$ denoting points, $\rho_{j}\geq 0$ weights) for velocity and energy respectively with $J_{v}$ and $J_{e}$ their respective sizes. We can approximate the integrations by
@@
The computational cost of the above summations is proportional to the sizes of the full quadrature rules ($J_{v}$ and $J_{e}$) since the integrands must be evaluated at every point of the corresponding quadrature rule. In addition, since the integrands depend on the state variables they must be updated at every stage of the time integration.
To remove the dependence of the computational cost on the full sizes $J_{v}$ and $J_{e}$ we employ a hyperreduction method called the empirical quadrature procedure (EQP) [57, 58]. EQP is used to generate quadrature rules of reduced size that can approximate the evaluation of the force functions with reduced computational cost while maintaining a reasonable level of accuracy. More specifically, EQP seeks to construct new quadrature rules $\{x^{v}_{j},\tilde{\rho}^{v}_{j}\}_{j=0}^{J_{v}-1}$ and $\{x^{e}_{j},\tilde{\rho}^{e}_{j}\}_{j=0}^{J_{e}-1}$ where only $\hat{J}_{v}\ll J_{v}$ and $\hat{J}_{e}\ll J_{e}$ of the new weights $\tilde{\rho}^{v}_{j}$ and $\tilde{\rho}^{e}_{j}$ are nonzero. In this way computing the sums in (3.13) requires the evaluation of the integrands at only $\hat{J}_{v}$ and $\hat{J}_{e}$ sampled quadrature points respectively, which are the points corresponding to nonzero weights.
Note that EQP does not attempt to identify new quadrature points; rather, it samples a subset of the original quadrature points with appropriately modified weights. In that sense EQP can be thought of as a quadrature analog of interpolation hyperreduction methods based on gappy POD [78]. The problem of constructing the reduced quadrature rules is posed as a linear optimization problem [57]. To outline the employed procedure we use the construction of a reduced quadrature rule for the approximation of $\bm{F}_{v}^{\psi}$ as an example. The construction of a reduced rule for $\bm{F}_{e}^{\psi}$ can be performed analogously.
Starting with the full quadrature rule $\{x^{v}_{j},\rho^{v}_{j}\}_{j=0}^{J_{v}-1}$ we want to find weights $\{\tilde{\rho}^{v}_{j}\}_{j=0}^{J_{v}-1}$ that minimize the $\ell^{1}$ norm $\sum_{j=0}^{J_{v}-1}\tilde{\rho}^{v}_{j}$ subject to the nonnegativity constraints $\tilde{\rho}^{v}_{j}\geq 0$ for all $j\in\{0,\ldots,J_{v}-1\}$ and a total of $N_{c}^{v}\ll J_{v}$ accuracy constraints
@@
where $0\leq k<N_{t}$ indexes the time snapshots $\tilde{w}(t_{k})$ used to form the constraints, $0\leq i<N_{v}$ indexes the components of vector $\bm{g}_{v}^{\psi}$ and $0\leq s<N_{c}^{v}$ is the overall index of the accuracy constraints with associated error thresholds $\epsilon_{s}>0$. In this work we set $N_{c}^{v}=n_{v}N_{t}$ which corresponds to $n_{v}$ constraints (one for each component of $\bm{g}_{v}^{\psi}$) per included snapshot of the state variables $\tilde{w}(t_{k})$. In addition to selecting the snapshots $\tilde{w}(t_{k})$ that will be used to form the accuracy constraints, selecting appropriate values for the error parameters $\epsilon_{s}$ is also important in constructing an effective reduced quadrature rule. In practice one seeks to strike a balance between sparsity (low number of nonzero weights, $\hat{J}_{v}$) and accuracy (low error thresholds $\epsilon_{s}$).
The linear optimization problem of forming the reduced quadrature rule can be recast as a nonnegative least squares (NNLS) problem [59, 79]. Details on the steps we take to form the NNLS problem and solve it are given in A.
3.2 Basic EQP hyperreduction
In this section we employ the EQP method as outlined above to derive a hyperreduced version of model (3.5). In particular, we use the reduced quadrature rules $\{x^{v}_{j},\tilde{\rho}^{v}_{j}\}_{j=0}^{J_{v}-1}$ and $\{x^{e}_{j},\tilde{\rho}^{e}_{j}\}_{j=0}^{J_{e}-1}$ to define the respective reduced force functions $\hat{\bm{F}}_{v}^{\psi}:\mathbb{R}^{n}\to\mathbb{R}^{n_{v}}$ and $\hat{\bm{F}}_{e}^{\psi}:\mathbb{R}^{n}\times\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{e}}$
@@ 3.2 Basic EQP hyperreduction
and semidiscrete hyperreduced model
@@
The evaluation of the reduced force functions requires only $\hat{J}_{v}\ll J_{v}$ and $\hat{J}_{e}\ll J_{e}$ evaluations of the respective integrands. As a result, it requires knowledge only of the degrees of freedom of $\tilde{w}$ sampled by the reduced quadrature rules. To make this reduction in computational cost explicit in our notation, the reduced force functions $\hat{\bm{F}}_{v}^{\psi}$ and $\hat{\bm{F}}_{e}^{\psi}$ are written as functions of the reduced state $\hat{\bm{w}}$ instead of the lifted state $\tilde{\bm{w}}$. For the evolution of $\hat{\bm{x}}$ in (3.15) the quantities $\bm{\Phi}_{x}^{\top}\bm{v}_{os}\in\mathbb{R}^{n_{x}}$ and $\bm{\Phi}_{x}^{\top}\bm{\Phi}_{v}\in\mathbb{R}^{n_{x}\times n_{v}}$ can be precomputed, leading to an evolution equation that depends on reduced sizes only.
@@ 3.3 Temporal discretization
Applying the RK2A scheme to the hyperreduced system (3.15) we arrive at the discrete hyperreduced system
@@
with $\bar{\hat{\bm{v}}}_{k+1/2}=(\hat{\bm{v}}_{k}+\hat{\bm{v}}_{k+1})/2$. In the above, the timestep control algorithm is enforced based on the lifted state vector $\tilde{\bm{w}}$, using again only the degrees of freedom sampled by the reduced quadrature rules. This generally leads to timesteps that are different from those used in the full order model where the timestep control algorithm is enforced using the full state vector $\bm{w}$.
@@ 3.4 Time windowing
For unsteady problems dominated by advection like the Euler equations considered in the present work, the dimension of the linear subspaces required to capture the relevant flow features generally grows with the simulation time. This is both because new dynamical behavior can emerge as the final simulation time is increased, and because dynamically similar flow features can appear in different regions of the spatial domain. As a result, attempting to build a single reduced model for the whole simulation time horizon may require a large reduced basis dimension to achieve the desired degree of accuracy. In turn, this can potentially eliminate any acceleration of the simulation offered by the employed reduction method.
To address this issue we use the method of time windowing [31, 32, 33, 34]. With time windowing, we divide the considered simulation time interval $t_{0},t_{f}]$ into $N_{w}$ subintervals (windows) $T_{i},T_{i+1}]$, $0\leq i<N_{w}$ and build a reduced model for each time window. More specifically, for each window we collect the solution snapshots that correspond to the prescribed time interval and use them to build reduced basis matrices and reduced quadrature rules that are local to the chosen time window. This allows us to build reduced bases and quadrature rules that are of modest size and offer high accuracy in approximating the full solution for each time subinterval.
@@ 3.5 Simulation stages
The overall simulation process consists of two main stages: the offline and online stages. In the offline stage we perform numerical simulations using the full model (2.8) and collect the desired solution snapshots. Using those snapshots we derive the reduced basis matrices and reduced quadrature rules. When multiple time windows are used this process is carried out for each time window independently as described earlier. These steps complete the construction of the hyperreduced model.
In the online stage we use the derived hyperreduced model to perform numerical simulations that depend on reduced sizes only, thereby offering a reduction in computation cost. In the present work we focus on reproductive simulation runs, where the derived reduced model is used to reproduce simulation data that has been used for its training. This is in contrast to predictive simulations, where out of sample initial conditions are used for testing the reduced model.
@@ 4 Energy conservative EQP
In this section we develop a modified version of the basic EQP method outlined earlier that enforces the additional constraint of total energy conservation in the discrete hyperreduced model. To do so, we use the semidiscrete system (3.7).
As before the goal is to derive reduced force functions $\hat{\bm{F}}_{v}^{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{n_{v}}$ and $\hat{\bm{F}}_{e}^{\phi}:\mathbb{R}^{n}\times\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{e}}$ to approximate the force function evaluations
@@
with reduced computational cost. With the reduced force functions in place, we derive the semidiscrete hyperreduced model
@@
and apply the RK2A scheme to reach the discrete hyperreduced system
@@
In the following sections we define the reduced force functions $\hat{\bm{F}}_{v}^{\phi}$ and $\hat{\bm{F}}_{e}^{\phi}$. We begin by identifying sufficient conditions for total energy conservation in the strong sense. Then we describe the implementation of the resulting hyperreduction scheme used to incorporate these conditions.

### v2

@@ 2.2 Temporal discretization
with $\bar{\bm{v}}_{k+1/2}=(\bm{v}_{k}+\bm{v}_{k+1})/2$. To ensure stability of the explicit timestepping scheme we also employ the timestep control algorithm described in [57].
@@ 3 Reduced model
The FEM coefficient vectors of the reduced basis functions are represented by the orthonormal basis matrices $\bm{\Phi}_{v}\in\mathbb{R}^{N_{v}\times n_{v}}$, $\bm{\Phi}_{e}\in\mathbb{R}^{N_{e}\times n_{e}}$ and $\bm{\Phi}_{x}\in\mathbb{R}^{N_{v}\times n_{x}}$, where each column holds the FEM coefficients of one basis function. We compute the reduced basis matrices from simulation data of the full model using the POD method [9]. To decide how many POD basis functions to use to form each basis matrix we employ the POD energy ratio
@@
where $\sigma_{i}\geq 0$ denotes the $i$th largest singular value of the employed data matrix consisting of $N_{snap}$ solution snapshots, $e_{\sigma}\geq 0$ a chosen threshold and $n_{k}$ the size of the corresponding reduced state vector (velocity $v$, energy $e$ or position $x$). Using the introduced reduced basis matrices we rewrite (3.1) in terms of FEM coefficient vectors
@@
with the appropriate initial condition at $t=0$. The above corresponds to using the reduced basis functions as the trial functions and the FEM basis functions as the test functions. Next we present the derivation of two alternative reduced systems of equations for the reduced state vector $\hat{\bm{w}}=(\hat{\bm{v}},\hat{\bm{e}},\hat{\bm{x}})\in\mathbb{R}^{n}$, $n=n_{v}+n_{e}+n_{x}$. The reason for introducing two alternative systems is that both systems will be employed by the hyper-reduction methods introduced in the following sections.
Starting from (3.4), the first reduced system is derived by inverting the mass matrices and transpose multiplying by the reduced basis matrices
@@
Although the size of the reduced systems (3.5) and (3.7) has been successfully reduced to $n\ll N$, the nonlinearity of the force functions in the right hand side of the equations means that they still depend on the full state vector $\tilde{\bm{w}}\in\mathbb{R}^{N}$. As a result, evaluating the force functions requires lifting the reduced state $\hat{\bm{w}}\in\mathbb{R}^{n}$ to $\tilde{\bm{w}}$ at every timestep of the integration. The cost of this lifting operation dominates the computational cost of the simulation and prevents its acceleration. To address this issue, a hyper-reduction method is employed to derive a hyper-reduced model that can be simulated without requiring lifting to the full state vector $\tilde{\bm{w}}$.
In this work we employ the EQP hyper-reduction method, which uses reduced quadrature rules to evaluate the nonlinear force functions with reduced cost [50, 51]. Quadrature methods such as EQP are naturally well suited to FEM spatial discretization procedures, such as the one employed in this work. In that respect, the present work can be viewed as a continuation of [22, 23], where interpolation methods were used to derive hyper-reduced Lagrangian hydrodynamics models.
@@ 3.1 Empirical quadrature procedure
In this section we describe the EQP hyper-reduction method using the model equations (3.5). The computation of the force vectors $\bm{F}_{v}^{\psi}$ and $\bm{F}_{e}^{\psi}$ of (3.5) requires the spatial integrations
@@
with the integrands $\bm{g}_{v}^{\psi}$ and $\bm{g}_{e}^{\psi}$ taking values in $\mathbb{R}^{n_{v}}$ and $\mathbb{R}^{n_{e}}$ respectively, with components
@@
These integrations can be approximated using numerical quadrature rules employing the underlying FEM basis functions [57, 60]. Let $\{x^{v}_{j},\rho^{v}_{j}\}_{j=0}^{J_{v}-1}$ and $\{x^{e}_{j},\rho^{e}_{j}\}_{j=0}^{J_{e}-1}$ denote the full quadrature rules ($x_{j}$ denoting points, $\rho_{j}\geq 0$ weights) for velocity and energy, with $J_{v}$ and $J_{e}$ their respective sizes. We can approximate the integrations by
@@
The computational cost of the above summations is proportional to the sizes of the full quadrature rules ($J_{v}$ and $J_{e}$), since the integrands must be evaluated at every point of the corresponding quadrature rule. In addition, since the integrands depend on the state variables they must be updated at every stage of the time integration.
To remove the dependence of the computational cost on the full sizes $J_{v}$ and $J_{e}$ we employ the EQP hyper-reduction method [50, 51]. EQP is used to generate quadrature rules of reduced size that can approximate the evaluation of the force functions with reduced computational cost while maintaining a reasonable level of accuracy. More specifically, EQP seeks to construct new quadrature rules $\{x^{v}_{j},\tilde{\rho}^{v}_{j}\}_{j=0}^{J_{v}-1}$ and $\{x^{e}_{j},\tilde{\rho}^{e}_{j}\}_{j=0}^{J_{e}-1}$ where only $\hat{J}_{v}\ll J_{v}$ and $\hat{J}_{e}\ll J_{e}$ of the new weights $\tilde{\rho}^{v}_{j}$ and $\tilde{\rho}^{e}_{j}$ are nonzero. In this way, computing the sums in (3.13) requires the evaluation of the integrands at only $\hat{J}_{v}$ and $\hat{J}_{e}$ sampled quadrature points respectively, which are the points corresponding to nonzero weights.
Note that EQP does not attempt to identify new quadrature points; rather, it samples a subset of the original quadrature points with appropriately modified weights. In that sense EQP can be thought of as a quadrature analog of interpolation hyper-reduction methods based on gappy POD [61]. The problem of constructing the reduced quadrature rules is posed as a linear optimization problem [50]. To outline the employed procedure we use the construction of a reduced quadrature rule for the approximation of $\bm{F}_{v}^{\psi}$ as an example. The construction of a reduced rule for $\bm{F}_{e}^{\psi}$ can be performed analogously.
Starting with the full quadrature rule $\{x^{v}_{j},\rho^{v}_{j}\}_{j=0}^{J_{v}-1}$ we want to find weights $\{\tilde{\rho}^{v}_{j}\}_{j=0}^{J_{v}-1}$ that minimize the 1-norm $\sum_{j=0}^{J_{v}-1}\tilde{\rho}^{v}_{j}$ subject to the nonnegativity constraints $\tilde{\rho}^{v}_{j}\geq 0$, $j\in\{0,\ldots,J_{v}-1\}$, and a total of $N_{c}^{v}\ll J_{v}$ accuracy constraints
@@
where $0\leq k<N_{t}$ indexes the time snapshots $\tilde{w}(t_{k})$ used to form the constraints, $0\leq i<n_{v}$ indexes the components of vector $\bm{g}_{v}^{\psi}$, and $0\leq s<N_{c}^{v}$ is the overall index of the accuracy constraints with associated error thresholds $\epsilon_{s}>0$. This implies a total of $N_{c}^{v}=n_{v}N_{t}$ constraints, corresponding to $n_{v}$ constraints (one for each component of $\bm{g}_{v}^{\psi}$) per included snapshot of the state variables $\tilde{w}(t_{k})$. In addition to selecting the snapshots $\tilde{w}(t_{k})$ that will be used to form the accuracy constraints, selecting appropriate values for the error thresholds $\epsilon_{s}$ is also important in constructing an effective reduced quadrature rule. In practice one seeks to strike a balance between sparsity (low number of nonzero weights, $\hat{J}_{v}$) and accuracy (low error thresholds $\epsilon_{s}$).
The linear optimization problem of forming the reduced quadrature rule can be recast as an equivalent nonnegative least squares (NNLS) problem [52, 62]. Details on the steps we take to form the NNLS problem and solve it are given in A.
3.2 Basic EQP hyper-reduction
In this section we employ the EQP method as outlined above to derive a hyper-reduced version of model (3.5). In particular, we use the reduced quadrature rules $\{x^{v}_{j},\tilde{\rho}^{v}_{j}\}_{j=0}^{J_{v}-1}$ and $\{x^{e}_{j},\tilde{\rho}^{e}_{j}\}_{j=0}^{J_{e}-1}$ to define the respective reduced force functions $\hat{\bm{F}}_{v}^{\psi}:\mathbb{R}^{n}\to\mathbb{R}^{n_{v}}$ and $\hat{\bm{F}}_{e}^{\psi}:\mathbb{R}^{n}\times\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{e}}$
@@ 3.2 Basic EQP hyper-reduction
and semidiscrete hyper-reduced model
@@
The evaluation of the reduced force functions requires only $\hat{J}_{v}\ll J_{v}$ and $\hat{J}_{e}\ll J_{e}$ evaluations of the respective integrands. As a result, it requires knowledge only of the degrees of freedom of $\tilde{w}$ sampled by the reduced quadrature rules. To make this reduction in computational cost explicit in our notation, the reduced force functions $\hat{\bm{F}}_{v}^{\psi}$ and $\hat{\bm{F}}_{e}^{\psi}$ are written as functions of the reduced state $\hat{\bm{w}}$, instead of the lifted state $\tilde{\bm{w}}$. For the evolution of $\hat{\bm{x}}$ in (3.15), the quantities $\bm{\Phi}_{x}^{\top}\bm{v}_{os}\in\mathbb{R}^{n_{x}}$ and $\bm{\Phi}_{x}^{\top}\bm{\Phi}_{v}\in\mathbb{R}^{n_{x}\times n_{v}}$ can be precomputed, leading to an evolution equation that depends on reduced sizes only.
@@ 3.3 Temporal discretization
Applying the RK2A scheme to the hyper-reduced system (3.15), we arrive at the discrete hyper-reduced system
@@
with $\bar{\hat{\bm{v}}}_{k+1/2}=(\hat{\bm{v}}_{k}+\hat{\bm{v}}_{k+1})/2$. In the above, the timestep control algorithm is enforced based on the lifted state vector $\tilde{\bm{w}}$, using again only the degrees of freedom sampled by the reduced quadrature rules. In general, this leads to timesteps that are different from those used in the full order model, where the timestep control algorithm is enforced using the full state vector $\bm{w}$.
@@ 3.4 Time windowing
For unsteady problems dominated by advection, like the Euler equations considered in the present work, the dimension of the linear subspaces required to capture the relevant flow features generally grows with the simulation time. This is both because new dynamical behavior can emerge as the final simulation time is increased, and because dynamically similar flow features can appear in different regions of the spatial domain. As a result, attempting to build a single reduced model for the whole simulation time horizon may require a large reduced basis dimension to achieve the desired degree of accuracy. In turn, this can potentially eliminate any acceleration of the simulation offered by the employed reduction method.
To address this issue we use the method of time windowing [20, 21, 22, 23]. With time windowing, we divide the considered simulation time interval $t_{0},t_{f}]$ into $N_{w}$ subintervals (windows) $t_{i},t_{i+1}]$, $0\leq i<N_{w}$, and build a reduced model for each time window. More specifically, for each window we collect the solution snapshots that correspond to the prescribed time interval and use them to build reduced basis matrices and reduced quadrature rules that are local to the chosen time interval. This allows us to build reduced bases and quadrature rules that are of modest size and offer high accuracy in approximating the full solution for each time subinterval.
@@ 3.5 Simulation stages
The overall simulation process consists of two main stages: the offline and online stages. In the offline stage we perform numerical simulations using the full model (2.8) and collect the desired solution snapshots. Using those snapshots, we derive the reduced basis matrices and reduced quadrature rules. When multiple time windows are used, this process is carried out for each time window independently as described earlier. These steps complete the construction of the hyper-reduced model.
In the online stage we use the derived hyper-reduced model to perform numerical simulations that depend on reduced sizes only, thereby offering a reduction in computational cost. In the present work we focus on reproductive simulations, where the derived reduced model is used to reproduce simulation data that has been used for its training. This is in contrast to predictive simulations, where the reduced model is tested in predicting previously unseen simulation data.
@@ 4 Energy conservative EQP
In this section we develop a modified version of the basic EQP method that enforces the additional constraint of total energy conservation in the discrete hyper-reduced model. To do so, we use the semidiscrete system (3.7). Our goal is to derive reduced force functions $\hat{\bm{F}}_{v}^{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{n_{v}}$ and $\hat{\bm{F}}_{e}^{\phi}:\mathbb{R}^{n}\times\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{e}}$ to approximate the force function evaluations
@@
with reduced computational cost. Using the reduced force functions, we can then derive the semidiscrete hyper-reduced model
@@
and apply the RK2A scheme to reach the discrete hyper-reduced system
@@
In the following sections we define the reduced force functions $\hat{\bm{F}}_{v}^{\phi}$ and $\hat{\bm{F}}_{e}^{\phi}$. We begin by identifying sufficient conditions for total energy conservation in the strong sense. Then we describe the implementation of the resulting hyper-reduction scheme used to incorporate these conditions.